A uniform magnetic field B=B*t(B* is a constant) fills a cylindrical volume of radius R,then the potential difference in the conducting rod PQ acting as a chord in the face of cylinder due to electrostatic field is:

Grade:

1 Answers

meghala
6 years ago
E=-Dphi/dt
=AdB/dt
where A=area of the triangle sandwiched between the ends of the rods and the radii connecting the centre with ends 
area=1/2*base*altitude
E=-db/dt(l/2)rootR^2-(l/2)^2